# Precision gravity constraints on large dark sectors

### v2

@@ 1 Introduction
Einstein’s general relativity (GR) continues to be the most successful classical description of gravity, having passed all experimental tests from the Solar System to binary pulsars with remarkable accuracy. However, as is by now well recognized, GR should be regarded as a low-energy effective field theory (EFT), valid below the Planck scale $M_{\rm Pl}$, where quantum effects become non-negligible Donoghue (1994a). Within this EFT framework, quantum loops of matter and gravitons induce calculable corrections to the Newtonian potential, with leading terms scaling as $1/r^{3}$ at large distances Bjerrum-Bohr et al. (2003); Holstein (2023); Faller (2008). The formalism of covariant nonlocal form factors, developed through the Barvinsky–Vilkovisky heat-kernel expansion, provides a powerful language to encode such quantum corrections Codello and Jain (2015, 2016).
Historically, the search for modifications to Newton’s law at short distances has been motivated by a variety of considerations, from the possibility of new light scalar mediators and axionlike particles, to the presence of large extra dimensions Adelberger et al. (2007); Kapner et al. (2007); Hoyle et al. (2004). Precision “fifth–force” experiments, such as torsion-balance measurements pioneered by the Eöt-Wash group, now probe gravitational interactions down to tens of microns Tan et al. (2020). These experiments are typically interpreted in terms of Yukawa-type deviations from the inverse-square law, parametrized by a strength $\alpha$ and range $\lambda$ Tan et al. (2020); Adelberger et al. (2007). The resulting exclusion curves are among the strongest probes of new macroscopic forces, complementing astrophysical and cosmological constraints.
In the context of hidden or dark sectors, quantum gravitational corrections acquire a qualitatively new role. A single new scalar, fermion, or vector state contributes negligibly to deviations from Newton’s law. However, a sector containing a large multiplicity $N$ of light degrees of freedom can lead to cumulative loop corrections that become appreciable. Massless dark particles (e.g. “dark photons”) generate enhanced $1/r^{3}$ power-law corrections, while massive states contribute Yukawa-suppressed tails whose collective strength grows with $N$ Burns and Pilaftsis (2014); Fröb (2016). This mechanism links the long-standing program of precision gravity tests with the particle-physics motivation to explore large hidden sectors, such as those emerging in dark QCD-like theories or string compactifications.
@@ 2.2 Effective Action and Nonlocal Form Factors
Quantum corrections to gravity can be systematically treated within the one-loop effective action, viewing general relativity as a low-energy EFT. For a matter field of mass $m$ and spin $s$ one writes Codello and Jain (2016, 2015)
@@
where $\Delta_{s}$ is the kinetic operator appropriate to spin $s$. The nonlocal terms of interest appear as form factors multiplying curvature invariants. It is convenient and complete to organize the result in the quadratic curvature basis (see, e.g., Barvinsky and Vilkovisky (1990); Donoghue (1994b); Donoghue and Holstein (2017); Burns and Pilaftsis (2015); Fröb et al. (2015)):
@@
$\Gamma^{(1)}_{(1)}\;=\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\mu}{}^{\nu}+m^{2}\big)\;-\;\mathrm{Tr}\ln\!\big(\Delta_{\rm gh}+m^{2}\big)\;+\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\rm gf}+m^{2}\big),$
@@
with the same nonlocal shape for both channels, $F_{i,1}(\tau)=c_{i,1}\,[\,1-\tau\ln(1+1/\tau)\,]$, but with spin-1 coefficients $c_{i,1}$ that include the ghost/gauge-fixing contributions. The static inverse–square–law combination probed by our fits is $F^{\rm(ISL)}_{1}(\tau)=F_{2,1}(\tau)-\tfrac{1}{2}F_{1,1}(\tau)$, so the net coefficient is $c_{1}\equiv c_{2,1}-\tfrac{1}{2}c_{1,1}.$ Including the Faddeev–Popov sector yields a negative value for this combination, $c_{1}=-1/(96\pi^{2})$; see, e.g., Barvinsky and Vilkovisky (1990); Burns and Pilaftsis (2015); Fröb et al. (2015) for explicit derivations in the covariant form-factor language.
@@
where $N_{s}$ counts the number of fields of spin $s$. If a non-minimal scalar coupling $\tfrac{1}{2}\xi R\phi^{2}$ is present, the trace (spin-0) channel carries the familiar $(1-6\xi)^{2}$ weight Barvinsky and Vilkovisky (1990); Burns and Pilaftsis (2015); in practice one may track this by $c_{0}\!\to c_{0}(1-6\xi)^{2}$ in (7).
In the massless limit ($\tau\to\infty$), $F_{i,s}(\tau)\simeq c_{i,s}(\ln\tau-2)$, reproducing the standard running $\ln(-\Box/\mu^{2})$ Barvinsky and Vilkovisky (1990); Donoghue (1994b). For $|\tau|\ll 1$ (momenta well below $m$),
$F_{i,s}(\tau)=c_{i,s}\Big(\frac{\tau}{6}-\frac{\tau^{2}}{60}+\cdots\Big),$ (8)
which yields the local derivative series generating the universal long-distance $1/r^{3}$ tail in the static potential. At intermediate distances the same physics is conveniently encoded by a Yukawa-envelope fit Fröb et al. (2015); Burns and Pilaftsis (2015). In this way, our mapping from ISL data to bounds on the cumulative contribution of hidden fields naturally accounts for both curvature–squared operators and their nonlocal running.
@@ 2.3 Running and Threshold Effects in $G(q^{2})$
where $\mu$ is a renormalization scale. This result is explicitly derived and discussed in detail later in Section 3. This universal logarithmic dependence is responsible for long-distance, nonlocal corrections to Newton’s law Donoghue (1994a); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002).
@@ 2.4 Correction to the Newtonian Potential
As outlined above, the corrections to the Newtonian potential are obtained by integrating the nonlocal form factor above for all possible particle spins $s$. In the case of the massless fields, such corrections to the gravitational potential induce long range effects of the form $O(1/r^{3})$ Donoghue (1994a); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002); Faller (2008); Holstein (2023):
@@
where  Bjerrum-Bohr et al. (2003)
@@
where $N_{f,s,v}$ count Dirac fermions, real scalars, and vectors. For massive fields, corrections are exponentially suppressed beyond the Compton wavelength and interpolate between short-distance (massless) and long-distance (suppressed) regimes. Explicit analytic and numerical forms are given in Fröb (2016); Burns and Pilaftsis (2014), including Bessel and Struve functions.
@@ 3 Massless Limit Consistency
A crucial consistency check for the quantum corrections derived in the preceding sections is that, in the limit of vanishing mass, the expressions for massive fields smoothly reproduce the well-known massless results both in momentum-space running and in the real-space potential. This property is explicitly realized in the referenced calculations Fröb (2016); Burns and Pilaftsis (2014).
@@ 3.1 Momentum-Space Running
For a field of mass $m$ and spin $s$, the nonlocal correction to Newton’s constant in momentum space is obtained via integrating the quantum corrections given by $F_{s}$ in Eq. LABEL:eq:quantum_corr to obtain the form  Fröb (2016):
@@
where $\mu$ has replaced the mass $m$ and defines the renormalization scale in the massless case. This matches the standard massless result, where the running of $G$ is governed by a $q^{2}\ln(-q^{2})$ form factor Donoghue (1994a); Bjerrum-Bohr et al. (2003). In the limit $m\rightarrow 0$, the nonlocal form factor reduces to the universal massless expression $\sim q^{2}ln(-q^{2})$, in agreement with earlier EFT analyses of graviton and massless matter loops. The scale m in the massive case thus plays the role of a regulator, which is replaced by a renormalization scale $\mu$ in the massless theory.
@@ 3.2 Position-Space Potential
Exponential/Yukawa: $e^{-mr}/r^{3}$ Nonanalytic $1/r^{3}$ tail Fröb (2016); Burns and Pilaftsis (2014)
Bessel/Struve function form Expands to $1/r^{3}$ Burns and Pilaftsis (2014)
@@
so that the contribution is short-ranged and becomes negligible beyond the particle’s Compton wavelength. More detailed treatments Burns and Pilaftsis (2014) show that the exact expression can be written in terms of special functions (e.g., Bessel and Struve), but the essential behavior is well captured by the exponential factor. Both Burns and Pilaftsis (2014); Fröb (2016) express the full quantum-corrected potential $V(r)$ for massive loops using Bessel/Struve functions:
@@ 4 Constraints on Hidden or Dark Sectors with Many Degrees of Freedom
The full quantum-corrected Newtonian potential arises from the combined effect of all light and heavy degrees of freedom running in loops  Bjerrum-Bohr et al. (2003); Fröb (2016). Each particle species contributes a correction determined by its spin, mass, and multiplicity, with massless states generating long-range, nonanalytic terms and massive states producing short-range Yukawa-suppressed effects. In practice, the net modification is obtained by summing over all particle types, weighted by the spin-dependent coefficients introduced in the previous section. This leads to the general expression for a modified gravitational potential:
@@
Note that our reinterpretation counts any propagating degree of freedom lighter than the experimental lever arm. Concretely, a state contributes to the potential if its mass satisfies $m\lesssim r_{\star}^{-1}$. In confining theories, heavy resonances above this scale are integrated out and do not affect our fits, whereas light PNGBs, glueballs, or other composites couple universally to gravity and contribute on the same footing as elementary fields. The bounds therefore depend only on the infrared spectrum, not on UV compositeness.
@@
where the detailed normalization depends on the spin, couplings, and explicit computation of $\beta_{X}$ for the field(s) in question Fröb (2016); Burns and Pilaftsis (2014). However, the mapping is not reliable as $m\to 0$; we return to a detailed discussion of this massless limit in Sec. 5.
@@ 4.3 Constraints and Projected Sensitivity
For instance, identifying $\lambda=1/(2m_{X})$ for a species of mass $m_{X}$, one finds that experiments typically constrain $|\gamma|\lesssim 10^{-2}$$10^{-4}$ at micron length scales Burns and Pilaftsis (2014). If the collective enhancement from a large number of species $N$ or from very light masses $m_{X}$ pushes the predicted quantum correction above these bounds, the corresponding models can be ruled out or tightly constrained. The overall size and sign of the effect depend on the weighted sum of coefficients $c_{s}$ for the contributing fields, and can therefore be evaluated model by model using the master expressions for $U(r)$ derived above Fröb (2016); Bjerrum-Bohr et al. (2003). If the hidden sector also contains massless states (such as scalars or dark photons), the cumulative correction is instead of the form
@@
onto the experimental Yukawa form with parameters $(\gamma,\lambda)$. Published exclusion plots in the $(\gamma,\lambda)$ plane then translate directly into limits on the number $N_{X}$ and mass $m_{X}$ of new species Fröb (2016); Burns and Pilaftsis (2014), as we will see below.
@@ 5.2 Massless dark degrees of freedom: closed–form bound with numbers
Using the most precise published ISL power–law constraint (Eöt–Wash–style torsion balance), [^1] Tan et al. quote power–law bounds including $k=3$; for convenience we take their tabulated $k=3$ value at 68% CL, $|\beta_{3}|\lesssim 7.5\times 10^{-5}$ Tan et al. (2020). Older analyses (e.g. Hoyle et al. (2004)) give consistent though weaker values. At 95% CL, numbers inflate by a factor $\sim\!1.6$–2.

### v3

@@ 1 Introduction
Einstein’s general relativity (GR) continues to be the most successful classical description of gravity, having passed all experimental tests from the Solar System to binary pulsars with remarkable accuracy. However, as is by now well recognized, GR should be regarded as a low-energy effective field theory (EFT), valid below the Planck scale $M_{\rm Pl}$, where quantum effects become non-negligible Donoghue:1994dn. Within this EFT framework, quantum loops of matter and gravitons induce calculable corrections to the Newtonian potential, with leading terms scaling as $1/r^{3}$ at large distances Bjerrum-Bohr:2002gqz; Holstein:2023qcy; Faller:2007sy. The formalism of covariant nonlocal form factors, developed through the Barvinsky–Vilkovisky heat-kernel expansion, provides a powerful language to encode such quantum corrections Codello:2015oqa; Codello:2015mba.
Historically, the search for modifications to Newton’s law at short distances has been motivated by a variety of considerations, from the possibility of new light scalar mediators and axionlike particles, to the presence of large extra dimensions Adelberger:2006dh; Kapner:2006si; Hoyle:2004cw. Precision “fifth–force” experiments, such as torsion-balance measurements pioneered by the Eöt-Wash group, now probe gravitational interactions down to tens of microns Tan:2019vwu. These experiments are typically interpreted in terms of Yukawa-type deviations from the inverse-square law, parametrized by a strength $\alpha$ and range $\lambda$ Tan:2019vwu; Adelberger:2006dh. The resulting exclusion curves are among the strongest probes of new macroscopic forces, complementing astrophysical and cosmological constraints.
In the context of hidden or dark sectors, quantum gravitational corrections acquire a qualitatively new role. A single new scalar, fermion, or vector state contributes negligibly to deviations from Newton’s law. However, a sector containing a large multiplicity $N$ of light degrees of freedom can lead to cumulative loop corrections that become appreciable. Massless dark particles (e.g. “dark photons”) generate enhanced $1/r^{3}$ power-law corrections, while massive states contribute Yukawa-suppressed tails whose collective strength grows with $N$ Burns:2014bva; Frob:2016lbc. This mechanism links the long-standing program of precision gravity tests with the particle-physics motivation to explore large hidden sectors, such as those emerging in dark QCD-like theories or string compactifications.
@@ 2.2 Effective Action and Nonlocal Form Factors
Quantum corrections to gravity can be systematically treated within the one-loop effective action, viewing general relativity as a low-energy EFT. For a matter field of mass $m$ and spin $s$ one writes Codello:2015mba; Codello:2015oqa
@@
where $\Delta_{s}$ is the kinetic operator appropriate to spin $s$. The nonlocal terms of interest appear as form factors multiplying curvature invariants. It is convenient and complete to organize the result in the quadratic curvature basis (see, e.g., Barvinsky:1990PT1; Donoghue:1994Player; Donoghue:2017EFTreview; Burns:2014FormFactors; Frob:2015Nonlocal):
@@
$\Gamma^{(1)}_{(1)}\;=\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\mu}{}^{\nu}+m^{2}\big)\;-\;\mathrm{Tr}\ln\!\big(\Delta_{\rm gh}+m^{2}\big)\;+\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\rm gf}+m^{2}\big),$ (5)
@@
with the same nonlocal shape for both channels, $F_{i,1}(\tau)=c_{i,1}\,[\,1-\tau\ln(1+1/\tau)\,]$, but with spin-1 coefficients $c_{i,1}$ that include the ghost/gauge-fixing contributions. The static inverse–square–law combination probed by our fits is $F^{\rm(ISL)}_{1}(\tau)=F_{2,1}(\tau)-\tfrac{1}{2}F_{1,1}(\tau)$, so the net coefficient is $c_{1}\equiv c_{2,1}-\tfrac{1}{2}c_{1,1}.$ Including the Faddeev–Popov sector yields a negative value for this combination, $c_{1}=-1/(96\pi^{2})$; see, e.g., Barvinsky:1990PT1; Burns:2014FormFactors; Frob:2015Nonlocal for explicit derivations in the covariant form-factor language.
It is important to emphasize that this curvature–squared form–factor coefficient $c_{1}$ is not identical to the coefficient of the spin–1 contribution in the long–distance potential written in Eq. (I.1) of Ref. Burns:2014FormFactors, where only the physical vector degrees of freedom contribute (counted by $N_{1}$), with the ghost sector omitted from that parametrization. After translating between the physical field–counting basis used in Eq. (I.1) and the covariant form–factor basis used here (which automatically incorporates the ghost sector via Eq. (5)), both approaches yield the same long–distance $1/r^{3}$ correction to the Newtonian potential. The negative value (7) therefore reflects the standard background–field definition of the ISL coefficient for spin–1 matter.
@@
where $N_{s}$ counts the number of fields of spin $s$. If a non-minimal scalar coupling $\tfrac{1}{2}\xi R\phi^{2}$ is present, the trace (spin-0) channel carries the familiar $(1-6\xi)^{2}$ weight Barvinsky:1990PT1; Burns:2014FormFactors; in practice one may track this by $c_{0}\!\to c_{0}(1-6\xi)^{2}$ in (8).
The long–distance $1/r^{3}$ correction to the Newtonian potential originates from the nonanalytic part of the covariant form factors $F_{i,s}(\tau)$. In the massless or high–momentum regime ($\tau\to\infty$), the form factors take the well–known limit $F_{i,s}(\tau)\simeq c_{i,s}(\ln\tau-2)$, reproducing the $\ln(-\Box/\mu^{2})$ running characteristic of massless loops Barvinsky:1990PT1; Donoghue:1994Player. For massive fields with momenta well below threshold ($|\tau|\ll 1$), the same nonlocal structure admits the local curvature expansion
$F_{i,s}(\tau)=c_{i,s}\Big(\tfrac{\tau}{6}-\tfrac{\tau^{2}}{60}+\cdots\Big),$ (9)
which is the low–momentum expansion of the form factor that yields the universal $1/r^{3}$ tail upon resummation. At intermediate distances, this physics is conveniently interpolated by a Yukawa–envelope fit Frob:2015Nonlocal; Burns:2014FormFactors. Our mapping from ISL data to constraints on hidden fields therefore consistently incorporates both the curvature–squared operators arising at low momenta and their nonlocal running inherited from the full form factor.
@@ 2.3 Running and Threshold Effects in $G(q^{2})$
where $\mu$ is a renormalization scale. This result is explicitly derived and discussed in detail later in Section 3. This universal logarithmic dependence is responsible for long-distance, nonlocal corrections to Newton’s law Donoghue:1994dn; Bjerrum-Bohr:2002gqz; Khriplovich:2002bt.
@@ 2.4 Correction to the Newtonian Potential
As outlined above, the corrections to the Newtonian potential are obtained by integrating the nonlocal form factor above for all possible particle spins $s$. In the case of the massless fields, such corrections to the gravitational potential induce long range effects of the form $O(1/r^{3})$ Donoghue:1994dn; Bjerrum-Bohr:2002gqz; Khriplovich:2002bt; Faller:2007sy; Holstein:2023qcy:
@@
where  Bjerrum-Bohr:2002gqz
@@
where $N_{f,s,v}$ count Dirac fermions, real scalars, and vectors. For massive fields, corrections are exponentially suppressed beyond the Compton wavelength and interpolate between short-distance (massless) and long-distance (suppressed) regimes. Explicit analytic and numerical forms are given in Frob:2016lbc; Burns:2014bva, including Bessel and Struve functions.
@@ 3 Massless Limit Consistency
A crucial consistency check for the quantum corrections derived in the preceding sections is that, in the limit of vanishing mass, the expressions for massive fields smoothly reproduce the well-known massless results both in momentum-space running and in the real-space potential. This property is explicitly realized in the referenced calculations Frob:2016lbc; Burns:2014bva.
@@ 3.1 Momentum-Space Running
For a field of mass $m$ and spin $s$, the nonlocal correction to Newton’s constant in momentum space is obtained via integrating the quantum corrections given by $F_{s}$ in Eq. LABEL:eq:quantum_corr to obtain the form  Frob:2016lbc:
@@
where $\mu$ has replaced the mass $m$ and defines the renormalization scale in the massless case. This matches the standard massless result, where the running of $G$ is governed by a $q^{2}\ln(-q^{2})$ form factor Donoghue:1994dn; Bjerrum-Bohr:2002gqz. In the limit $m\rightarrow 0$, the nonlocal form factor reduces to the universal massless expression $\sim q^{2}ln(-q^{2})$, in agreement with earlier EFT analyses of graviton and massless matter loops. The scale m in the massive case thus plays the role of a regulator, which is replaced by a renormalization scale $\mu$ in the massless theory.
@@ 3.2 Position-Space Potential
Exponential/Yukawa: $e^{-mr}/r^{3}$ Nonanalytic $1/r^{3}$ tail Frob:2016lbc; Burns:2014bva
Bessel/Struve function form Expands to $1/r^{3}$ Burns:2014bva
@@
so that the contribution is short-ranged and becomes negligible beyond the particle’s Compton wavelength. More detailed treatments Burns:2014bva show that the exact expression can be written in terms of special functions (e.g., Bessel and Struve), but the essential behavior is well captured by the exponential factor. Both Burns:2014bva; Frob:2016lbc express the full quantum-corrected potential $V(r)$ for massive loops using Bessel/Struve functions:
@@ 4 Constraints on Hidden or Dark Sectors with Many Degrees of Freedom
The full quantum-corrected Newtonian potential arises from the combined effect of all light and heavy degrees of freedom running in loops  Bjerrum-Bohr:2002gqz; Frob:2016lbc. Each particle species contributes a correction determined by its spin, mass, and multiplicity, with massless states generating long-range, nonanalytic terms and massive states producing short-range Yukawa-suppressed effects. In practice, the net modification is obtained by summing over all particle types, weighted by the spin-dependent coefficients introduced in the previous section. This leads to the general expression for a modified gravitational potential:
@@
Note that our reinterpretation counts any propagating degree of freedom lighter than the experimental lever arm. Concretely, a state contributes to the potential if its mass satisfies $m\lesssim r_{\star}^{-1}$. In confining theories, heavy resonances above this scale are integrated out and do not affect our fits, whereas light PNGBs, glueballs, or other composites couple universally to gravity and contribute on the same footing as elementary fields. Our constraints apply to any light degree of freedom that behaves as an effectively pointlike scalar at the momentum transfers relevant for ISL experiments. For PNGBs, this condition is naturally satisfied because their masses can be parametrically smaller than their compositeness scale. For glueballs, this occurs only if the confinement scale of the hidden sector is sufficiently low (e.g. keV–MeV), so that the lightest glueball is both kinematically accessible and structureless at the probed momenta. In contrast, QCD glueballs are too heavy to satisfy this condition, and do not enter our bounds. The bounds therefore depend only on the infrared spectrum, not on UV compositeness.
@@
where the detailed normalization depends on the spin, couplings, and explicit computation of $\beta_{X}$ for the field(s) in question Frob:2016lbc; Burns:2014bva. However, the mapping is not reliable as $m\to 0$; we return to a detailed discussion of this massless limit in Sec. 5.
@@ 4.3 Constraints and Projected Sensitivity
For instance, identifying $\lambda=1/(2m_{X})$ for a species of mass $m_{X}$, one finds that experiments typically constrain $|\gamma|\lesssim 10^{-2}$$10^{-4}$ at micron length scales Burns:2014bva. If the collective enhancement from a large number of species $N$ or from very light masses $m_{X}$ pushes the predicted quantum correction above these bounds, the corresponding models can be ruled out or tightly constrained. The overall size and sign of the effect depend on the weighted sum of coefficients $c_{s}$ for the contributing fields, and can therefore be evaluated model by model using the master expressions for $U(r)$ derived above Frob:2016lbc; Bjerrum-Bohr:2002gqz. If the hidden sector also contains massless states (such as scalars or dark photons), the cumulative correction is instead of the form
@@
onto the experimental Yukawa form with parameters $(\gamma,\lambda)$. Published exclusion plots in the $(\gamma,\lambda)$ plane then translate directly into limits on the number $N_{X}$ and mass $m_{X}$ of new species Frob:2016lbc; Burns:2014bva, as we will see below.
@@ 5.2 Massless dark degrees of freedom: closed–form bound with numbers
Using the most precise published ISL power–law constraint (Eöt–Wash–style torsion balance), [^1] Tan et al. quote power–law bounds including $k=3$; for convenience we take their tabulated $k=3$ value at 68% CL, $|\beta_{3}|\lesssim 7.5\times 10^{-5}$ Tan:2019vwu. Older analyses (e.g. Hoyle:2004cw) give consistent though weaker values. At 95% CL, numbers inflate by a factor $\sim\!1.6$–2.